OpenStudy (anonymous):

The graph of which of the following will be parallel to the graph of 4x – 3y = –12? y =4/3x - 3/2 6x - 4y = -8 y = 3/4x + 1 4x = 2y = -12 (I think it's either the 3rd or 4th one, but i'm leaning more towards the last one)

OpenStudy (texaschic101):

Put your equation in y = mx + b form, and m is your slope 4x - 3y = -12 -3y = -4x - 12 y = 4/3x + 4 slope is 4/3 parallel line will have the same slope...so which equation has the same slope (4/3) ?

OpenStudy (anonymous):

The second one?

OpenStudy (anonymous):

Wait, no, the first one!

OpenStudy (texaschic101):

you got it...the first one ...good job :)